OpenStudy (anonymous):

Solve for the variable X-1 over 8 = 3 over 4

Parth (parthkohli):

Try cross multiplying.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4(x - 1) = 8\cdot 3 }\)

Parth (parthkohli):

Can you manage it now?

OpenStudy (anonymous):

no

Parth (parthkohli):

Try distributing using the distributive property - \(a(b \pm c ) = ab \pm ac\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 4x - 4 = 24}\) Isolate \(x\).

OpenStudy (anonymous):

X-1/8=3/4 4(X-1)=3*8 4X-4=24 4X=24+4 4X=28 X=28/4 X=7